OpenStudy (anonymous):

What is the slope and y-intercept of this line? y = x – 3

OpenStudy (anonymous):

Given two points (x1, y1) and (x2, y2), the formula for the slope of the straight line going through these two points is: slope formula: m = [y1 - y2] / [x1 - x2] ...where the subscripts merely indicate that you have a "first" point (whose coordinates are subscripted with a "1") and a "second" point (whose coordinates are subscripted with a "2"); that is, the subscripts indicate nothing more than the fact that you have two points to work with. Note that the point you pick as the "first" one is irrelevant; if you pick the other point to be "first", then you get the same value for the slope: m = (y2 - y1) / (x2 - x1) ADVERTISEMENT (If you're not sure that the two formulas above give exactly the same values, no matter the pair of points plugged into them, then pick some points and try them out. See what you get.) The formula for slope is sometimes referred to as "rise over run", because the fraction consists of the "rise" (the change in y, going up or down) divided by the "run" (the change in x, going from left to the right). If you've ever done roofing, built a staircase, graded landscaping, or installed gutters or outflow piping, you've probably encountered this "rise over run" concept. The point is that slope tells you how much y is changing for every so much that x is changing. Pictures can be helpful, so let's look at the line y = ( 2/3 )x – 4; we'll compute the slope, and draw the line. To find points from the line equation, we have to pick values for one of the variables, and then compute the corresponding value of the other variable. If, say, x = –3, then y = ( 2/3 )(–3) – 4 = –2 – 4 = –6, so the point (–3, –6) is on the line. If x = 0, then y = ( 2/3 )(0) – 4 = 0 – 4 = –4, so the point (0, –4) is on the line. Now that we have two points on the line, we can find the slope of that line from the slope formula: slope m = 2/3 Let's look at these two points on the graph: Copyright © Elizabeth Stapel 2000-2011 All Rights Reserved

OpenStudy (anonymous):

lol

OpenStudy (jojo4eva):

The slope is 1 and the y intercept is -3

OpenStudy (jojo4eva):

Got it?? @kpopaigoo

OpenStudy (anonymous):

i don't think she does

OpenStudy (anonymous):

excuse you @d&r actually i do i just left for a moment and thank you @jojo4eva

OpenStudy (jojo4eva):

@kpopaigoo your welcome hun anytime you need anything don't hesitate to tag me i got you :) and @d&r that was completely unnecessary